OpenStudy (anonymous):

Find the equation of the tangent to the curve y = -cotx at the point (pi/4, -1) Show any derivatives that you need to find when solving this problem

OpenStudy (anonymous):

I know that -cotx differentiates to cosec2x but idk how to progress further

OpenStudy (anonymous):

Develop an equation for a straight line using point-slope form: \[(y-y1) = m(x-x1)\] where (x1,y1) = (pi/4,-1) and m = y'(pi/4)

OpenStudy (anonymous):

How do i differentiate cosec^2x and (pi/4)?

OpenStudy (anonymous):

\[y \prime (x) = - csc^2(x)\] Solve for y'(pi/4)

OpenStudy (anonymous):

You dont differentiate again

OpenStudy (anonymous):

Ok so how do you differentiate (pi/4) though?

OpenStudy (anonymous):

The answer says 2 but idk how they got that

OpenStudy (anonymous):

You plug in pi/4 for x in the above equation then use for the straight line equation

OpenStudy (anonymous):

ok thanks, I got the final answer now

OpenStudy (anonymous):

y = 2x-2.57